# SparseVLM: Visual Token Sparsification for Efficient Vision-Language Model Inference

### v2

@@ 3.3 Visual Token Recycling
We progressively sparsify visual tokens in each layer in the decoder, which results in more discarded tokens at later stages. Despite less significant, the pruned vision tokens with relatively large values in ${\bm{P}}$ still contain certain information. To efficiently preserve more visual details with fewer tokens, we propose a token recycling strategy to aggregate and reconstruct tokens to be sparsified.
@@ 3.4 Theoretical Analysis of Computational Complexity
We consider the computation of multi-head attention and feed-forward network (FFN) modules in the FLOPs estimation. Assuming $N$ is the number of pruned tokens, $D$ is the hidden state size, which is the same as the intermediate size in FFN, the FLOPs for one transformer layer can be reduced by $6(N-C)D^{2}+2(N-C)^{2}D$. Besides, our framework introduces additional computational overhead for the sparsification step with the details provided in Appendix A.2. Thus, we estimate the FLOP savings as the reduction part minus the additional sparsification overhead computed as
$\begin{array}[]{ll}\underbrace{\scriptstyle{\textstyle\sum_{i}}6(N_{i}-C_{i})D^{2}+2(N_{i}-C_{i})^{2}D}_{\text{reduction part}}-\underbrace{\scriptstyle 2L_{t}L_{v}D-{\textstyle\sum_{i}}L_{t}^{i}L_{v}^{i}(1+\min(L_{t}^{i},L_{v}^{i}))-(6{L^{i}_{r}}^{2}+2L_{r}^{i})D-L_{r}^{i}}_{\text{overhead part}}\\ \approx-2L_{t}L_{v}D+\sum_{i}D(6DN_{i}(1-x)+N_{i}^{2}(2+2x^{2}-4x-6(\tau)^{2}))-{L^{i}_{t}}^{2}L_{v}^{i}\\ \approx-2L_{t}L_{v}D+\sum_{i}DN_{i}(6D+2N_{i})-{L^{i}_{t}}^{2}L_{v}^{i},\ \ i\in\{1,2,\ldots,\Omega\},\end{array}$ (12)
where $\Omega$ is the number of total layers, and $x=\tau\times\theta$ is a very small decimal that can be ignored.
@@ 4 Experiments
Table 1: Performance of SparseLLaVA under different vision token configurations. The vanilla number of vision tokens is $576$. The first line of each method is the raw accuracy of benchmarks, and the second line is the proportion relative to the upper limit. The last column is the average value.
@@
ToMe (ICLR23) 54.3 60.5 1563 72.4 65.2 68.0 52.1 17.4 88.4%
87.7% 93.5% 83.9% 84.3% 93.8% 86.6% 89.5% 87.9%
FastV (ECCV24) 52.7 61.2 1612 64.8 67.3 67.1 52.5 18.0 88.1%
85.1% 94.6% 86.6% 75.4% 96.8% 85.5% 90.2% 90.9%
SparseVLM 57.6 62.5 1721 83.6 69.1 75.6 56.1 18.8 95.8%
@@
ToMe (ICLR23) 52.4 53.3 1343 62.8 59.6 63.0 49.1 16.0 80.4%
84.7% 82.4% 72.1% 73.1% 85.8% 80.2% 84.4% 80.8%
FastV (ECCV24) 49.6 56.1 1490 59.6 60.2 61.8 50.6 17.1 81.9%
80.1% 86.7% 80.0% 69.4% 86.6% 78.7% 86.9% 86.4%
SparseVLM 56.0 60.0 1696 80.5 67.1 73.8 54.9 18.5 93.3%
@@
ToMe (ICLR23) 48.6 43.7 1138 52.5 50.0 57.1 45.3 14.0 70.2%
78.5% 67.5% 61.1% 61.1% 71.9% 72.7% 77.8% 70.7%
FastV (ECCV24) 46.1 48.0 1256 48.0 51.1 55.0 47.8 15.6 72.1%
74.5% 74.2% 67.5% 55.9% 73.5% 70.1% 82.1% 78.8%
SparseVLM 52.7 56.2 1505 75.1 62.2 68.2 51.8 17.7 86.9%
@@
In this section, we validate our method within various VLM architectures on comprehensive multimodal benchmarks to assess its effectiveness including image and video understanding tasks.
@@ 4.1 Image Understanding Tasks
Implementation Details. We verify the proposed SparseVLM on two popular VLM frameworks: LLaVA (Liu et al., 2024) and Mini-Gemini (Li et al., 2024b). LLaVA-1.5 employs CLIP-pretrained ViT-L as the visual tower, while Mini-Gemini (MGM) further introduces a LAION-pretrained ConvNeXt-L (Liu et al., 2022) for high-resolution refinement. For LLaVA-1.5-7/13B and Mini-Gemini, we follow the same inference setting as the original paper as it is publicly available111github.com/haotian-liu/LLaVA and github.com/dvlab-research/MGM..
Main Results. In Table 1, we present the performance of SparseLLaVA (LLaVA equipped with SparseVLM) on image understanding benchmarks. To intuitively assess the performance, we provide the results by percentage format for comparative analysis, and the accuracy of the vanilla model with the 100% upper limit. We set $3$ vision token count configurations ($192$, $128$, and $64$) to check the advantages of SparseVLM. When pruning from $576$ to $192$ tokens, the SparseLLaVA only decreases the average accuracy by $4.2\%$ without additional training and exceeds ToMe (Bolya et al., 2022) $7.4\%$. Furthermore, when only $64$ tokens are left, our method outperforms FastV (Chen et al., 2024b) by a significant margin of 14.8%, while ToMe performs worst due to its direct merging.
Figure 4 shows the performance of SparseMGM, and we visualize the results on POPE, TextVQA, and GQA. We find that our framework has an obvious advantage over FastV and ToMe. With the reduction of tokens, the gap between FastV and SparseVLM is increasing sharply. The reason is that, compared to FastV and ToMe, the text-aware strategy enables us to accurately locate visual tokens with more details, while the recycling of pruned tokens further reduces information loss.
Figure 4: Performance of MGM armed with SparseVLM on three multimodal benchmarks. The horizontal axis represents the remaining number of vision tokens, while the vertical axis means the accuracy after percentage normalization. FastV is included for comparison.
@@ 4.2 Video Understanding Tasks
Table 2: The results of Video-LLaVA with SparseVLM on video question answering task. The original number of video tokens is $2048$, while our experiment collectively prunes it down to $135$ tokens. FastV is included for comparison. The GPT-3.5 turbo is adopted for assistive evaluation.
Method TGIF MSVD MSRVTT ActivityNet Avg
Acc Score Acc Score Acc Score Acc Score Acc Score
Video-LLaVA 47.1 3.35 69.8 3.92 56.7 3.48 43.1 3.35 100.0% +0.00
FastV (ECCV24) 23.1 2.47 38.0 2.71 19.3 2.02 30.6 2.82 52.1% -1.02
49.0% -0.88 54.4% -1.21 34.0% -1.46 71.0% -0.53
SparseVLM 44.7 3.29 68.2 3.90 31.0 2.68 42.6 3.32 86.5% -0.17

### v3

@@ 3.3 Visual Token Recycling
We progressively sparsify visual tokens in each layer in the decoder, which results in more discarded tokens at later stages. Despite being less significant, the pruned visual tokens with relatively large values in ${\bm{P}}$ still contain certain information. To efficiently preserve more visual details with fewer tokens, we propose a token recycling strategy to aggregate and reconstruct tokens to be pruned.
@@ 3.4 Theoretical Analysis of Computational Complexity
We consider the computation of multi-head attention and feed-forward network (FFN) modules in the FLOPs estimation. Assuming $N$ is the number of pruned tokens, $D$ is the hidden state size, which is the same as the intermediate size in FFN, the FLOPs for one Transformer layer can be reduced by $6(N-C)D^{2}+2(N-C)^{2}D$. Besides, our partial step introduces minimal computation with the details provided in Appendix A.3. Thus, we estimate the FLOPs savings as the reduction part minus the additional overhead:
$\begin{array}[]{ll}\underbrace{\scriptstyle{\textstyle\sum_{i}}6(N_{i}-C_{i})D^{2}+2(N_{i}-C_{i})^{2}D}_{\text{reduction part}}-\\ \underbrace{\scriptstyle 2L_{t}L_{v}D-{\textstyle\sum_{i}}L_{t}^{i}L_{v}^{i}(1+\min(L_{t}^{i},L_{v}^{i}))-(6{L^{i}_{r}}^{2}+2L_{r}^{i})D-L_{r}^{i}}_{\text{overhead part}}\\ \approx\scriptstyle-2L_{t}L_{v}D+\sum_{i}D(6DN_{i}(1-x)+N_{i}^{2}(2+2x^{2}-4x-6(\tau)^{2}))-{L^{i}_{t}}^{2}L_{v}^{i}\\ \approx\scriptstyle-2L_{t}L_{v}D+\sum_{i}DN_{i}(6D+2N_{i})-{L^{i}_{t}}^{2}L_{v}^{i},\end{array}$ (11)
where $i\in\{1,2,\ldots,\Omega\}$ and $\Omega$ is the number of total layers, and $x=\tau\times\theta$ is a very small decimal that can be ignored.
@@ 4 Experiments
Table 1: Performance of SparseLLaVA under different vision token configurations. The vanilla number of vision tokens is $576$. The first line of each method is the raw accuracy of benchmarks, and the second line is the proportion relative to the upper limit.
@@
ToMe (ICLR23) 54.3 60.5 1563 72.4 65.2 53.1 52.1 27.9 88.9 $(\downarrow 11.1)$ 2.05 34.06
87.7% 93.5% 83.9% 84.3% 93.8% 88.1% 89.5% 90.3%
FastV (ECCV24) 52.6 61.0 1605 64.8 69.1 52.1 52.5 26.7 87.9 $(\downarrow 12.1)$ 2.11 34.87
85.0% 94.4% 86.1% 75.4% 99.4% 86.4% 90.1% 86.4%
PDrop (Oct.24) 57.1 63.2 1766 82.3 70.2 54.7 56.1 30.5 95.9 $(\downarrow 4.1)$ 2.03 36.74
92.2% 97.8% 94.7% 95.8% 101.0% 90.7% 96.2% 98.7%
@@
ToMe (ICLR23) 52.4 53.3 1343 62.8 59.6 50.9 49.1 27.2 81.9 $(\downarrow 18.1)$ 1.62 30.00
84.7% 82.4% 72.1% 73.1% 85.8% 84.4% 84.4% 88.0%
FastV (ECCV24) 49.6 56.1 1490 53.4 68.6 48.1 50.5 26.3 82.4 $(\downarrow 17.6)$ 1.70 30.70
80.1% 86.8% 79.9% 62.2% 98.7% 79.8% 86.6% 85.1%
PDrop (Oct.24) 56.0 61.1 1664 82.3 69.9 53.3 55.1 30.8 94.3 $(\downarrow 5.7)$ 1.62 37.77
90.5% 95.4% 89.3% 95.8% 100.6% 88.4% 94.5% 99.7%
@@
ToMe (ICLR23) 48.6 43.7 1138 52.5 50.0 44.0 45.3 24.1 71.1 $(\downarrow 28.9)$ 1.19 26.52
78.5% 67.5% 61.1% 61.1% 71.9% 73.0% 77.8% 78.0%
FastV (ECCV24) 46.1 47.2 1255 38.2 68.7 43.7 47.8 19.6 72.0 $(\downarrow 28.0)$ 1.29 27.30
74.5% 73.1% 67.3% 44.5% 98.8% 72.5% 82.0% 63.4%
PDrop (Oct.24) 41.9 33.3 1092 55.9 69.2 40.0 45.9 30.7 73.4 $(\downarrow 26.6)$ 1.18 43.41
67.7% 51.6% 58.6% 65.1% 99.6% 66.3% 78.7% 99.4%
@@
86.9% 93.0% 85.2% 90.2% 100.4% 86.6% 91.6% 80.6%
In this section, we validate our method within various vision-language architectures on comprehensive multimodal benchmarks, including image and video understanding tasks, to assess its generality, effectiveness, and efficiency.
@@ 4.1 Image Understanding Tasks
Implementation Details. We verify SparseVLM on three VLM frameworks: LLaVA (Liu et al., 2024), Mini-Gemini (MGM) (Li et al., 2024c), and Qwen2-VL (Bai et al., 2023). LLaVA-1.5 employs CLIP-pretrained ViT-L as the visual tower, MGM further introduces a LAION-pretrained ConvNeXt-L (Liu et al., 2022) for high-resolution refinement, while Qwen2-VL owns dynamic resolution encoder.
Main Results. In Table 1, we present the performance of SparseLLaVA (LLaVA equipped with SparseVLM) on image understanding benchmarks. To intuitively assess the performance, we provide the results by percentage format for comparative analysis, and the accuracy of the vanilla model with the 100% upper limit. We set $3$ vision token count configurations ($192$, $128$, and $64$) to check the advantages of SparseVLM comprehensively. When pruning from $576$ to $192$ tokens, the SparseLLaVA only decreases the average accuracy by $0.9\%$ without additional training and exceeds ToMe (Bolya et al., 2022) $10.2\%$. When only $64$ tokens are kept, our method outperforms FastV (Chen et al., 2024) by a significant margin of 17.3%, while ToMe performs worst due to its direct merging. Furthermore, we also compare the recent method PDrop (Xing et al., 2024) training-free version, which has lower FLOPs computation. However, our method outperforms it in accuracy and latency, which are the most crucial metrics for practical deployment.
Figure 4 visualizes the performance of SparseMGM on POPE, TextVQA, and GQA. We find that our framework has an obvious advantage over FastV and ToMe. With the reduction of tokens, the gap between FastV and SparseVLM is increasing sharply. The reason is that, compared to FastV and ToMe, the text-aware strategy enables us to accurately locate visual tokens with more details, while the recycling of pruned tokens further reduces information loss.
Table 2: Performance of SparseVLM on Qwen2-VL.
[…]
Dynamic 80.5 (1323) 86.4 (1311) 84.3 (1326) 83.7
600 79.6 86.5 80.3 82.1
500 78.8 86.3 79.0 81.4
400 79.0 85.8 77.1 80.7
We further investigate our efficacy on Qwen2-VL. In Table 2, when $54.5\%$ of vision tokens are removed, Qwen2-VL maintains an accuracy of $98.0\%$. Furthermore, for every 100 tokens pruned, the accuracy only drops by approximately $0.8\%$. This validates the effectiveness of our method at high resolutions and its compatibility with variable resolutions.
Figure 4: Performance of MGM w/ SparseVLM on three multimodal benchmarks. The horizontal axis represents the remaining number of vision tokens, while the vertical axis means the accuracy after percentage normalization.
@@ 4.2 Video Understanding Tasks
Table 3: The results of Video-LLaVA with SparseVLM on video question answering task. The original number of video tokens is $2048$, while our experiment collectively prunes it down to $198$ tokens. FastV (Chen et al., 2024) is included for comparison.
Method TGIF MSVD MSRVTT ActivityNet Avg.
Video-LLaVA 47.1 69.8 56.7 43.1 100.0%
100% 100% 100% 100%
FastV 23.1 38.0 19.3 30.6 52.1%
49.0% 54.4% 34.0% 71.0%
Ours 44.7 68.2 31.0 42.6 86.5%
94.9% 97.7% 54.7% 98.8%
Datasets. We test on four common video question answering benchmarks, TGIF-QA (Jang et al., 2017), MSVD-QA (Xu et al., 2017), MSRVTT-QA (Xu et al., 2017), and ActivityNet-QA (Yu et al., 2019), where video-question pairs are massively disproportional in length. We use the evaluation tool of Video-ChatGPT (Maaz et al., 2023).
Implementation Details. We directly apply our SparseVLM for Video-LLaVA (Lin et al., 2023), which is composed of several key components, including language bind encoder $f^{v}_{M}$(Zhu et al., 2023a) for extracting features from raw visual inputs (e.g., images or videos), a language decoder model $f_{L}$ such as Vicuna (Touvron et al., 2023), a visual projection layer $f_{P}$, and a word embedding layer $f_{T}$.
Main Results. In Table 3, we set the Video-LLaVA with $2048$ video tokens as our upper bound for an overall average accuracy of $100.0\%$ and a score of $+0.00$. To make a fair comparison, we both preserve $198$ vision tokens ($90.3\%$ pruning ratio) for FastV (Chen et al., 2024) and SparseVLM. It is clear that our approach consistently outperforms FastV across all benchmarks, both in accuracy (Acc.) and GPT evaluation score. SparseVideoLLaVA achieves a total average accuracy of $86.5\%$, a significant 34.4% higher than $52.1\%$ of FastV. (From the GPT score perspective, SparseVLM only loses $0.17$ points compared to $1.02$ points of FastV.) These improvements suggest that when handling video modality containing temporal features, SparseVLM continues to deliver strong performance, generating accurate responses to diverse questions while utilizing significantly fewer tokens. This achieves an effective trade-off between inference efficiency and model performance.
5 Analysis
5.1 Relevant Text Token Selection
Figure 5: The ablation study of text raters on LLaVA 7B.
We propose a selection mechanism to localize visually irrelevant text tokens to limit their negative effects in rating the significance of vision tokens. Here we conduct experiments to analyze the effects of the mechanism in Figure 5. Under the same number of vision tokens ($64$), we have 3 settings (using all tokens, only text tokens, and only text raters we select) with LLaVA (Liu et al., 2023a) to judge vision token candidates. In TextVQA (Singh et al., 2019), by building upon the text-aware manner, our mechanism improves the baseline (all tokens) by $0.8\%$, which validates that our extra selection is effective. Besides, we further outperform the vanilla text-aware method (only text tokens) by $\textbf{2.7}\%$ on POPE (Li et al., 2023b). The huge margin means POPE sparsification is quite sensitive to question prompts, and text guidance is necessary. In summary, text rater selection is general and improves the performance across scenarios.
5.2 Recycling of Pruned Tokens
Table 4: Ablation study on token reconstruction (TR). Experiments are conducted on GQA and POPE on LLaVA 7B.
Benchmark Tokens Avg.
64 96 128 192
GQA 52.2 55.2 58.1 59.4 56.2
+ TR 53.8 56.4 58.4 59.5 57.0
POPE 72.8 77.5 83.7 85.2 79.8
+ TR 77.5 81.9 85.0 85.3 82.4
5.3 Computational Efficiency
SparseVLM affords significant efficiency and storage gains for the inference process. We conduct a comparative analysis of CUDA time, and FLOPs on LLaVA-7B, and compare our method with the baseline method and FastV (Chen et al., 2024). As displayed in Table 1, we conduct an inference efficiency analysis on a single NVIDIA A100-80GB with identical lengths of text prompts and single-image inputs. Compared to the baseline model, SparseVLM achieves a significant reduction of $43.1\%$ in CUDA time and $62.8\%$ in FLOPs while keeping $96.7\%$ accuracy. Despite SparseVLM has a minimal overhead to calculate text raters and cluster-pruned vision tokens, it leads to fewer than FastV tokens with comparable accuracy. Additionally, SparseVLM saves $67\%$ cache memory compared to vanilla LLaVA (where 302.4MB is reduced to 100.8MB), while keeping $99.1\%$ accuracy. More efficiency visualization (e.g., efficiency on VideoLLaVA) can be found in the Appendix A.8.
5.4 Qualitative Visualization
As shown in Figure 6, we visualize SparseVLM on various VQA questions. From left to right, we visualize the results after we apply token pruning to different layers. As the number of layers increases, more tokens are pruned and the Region of Interest (ROI) is gradually refined. The model systematically reduces less relevant image information while retaining key tokens closely tied to the question. The visualization reveals that SparseVLM, although discarding some overall image details, effectively retains essential visual tokens. These preserved tokens encapsulate the features necessary for answering the question, focusing on more relevant visual regions through their interaction with the question. More cases are in the Appendix A.7.
6 Conclusion
This paper introduced a text-aware training-free token optimization approach called SparseVLM which significantly decreased the test-time computations of various VLMs. Unlike prior methods, SparseVLM optimized VLMs without introducing extra parameters and fine-tuning costs. We achieved a more compact visual representation by employing the rank of attention matrices to determine pruning ratios and by recycling the pruned tokens via the reconstruction mechanism to reduce the information loss. Experiments demonstrated that e.g. the LLaVA when equipped with SparseVLM achieved 37.0% reduction in latency with a compression ratio of 77.8% while maintaining 97% of the original accuracy. Moreover, our method exceeded FastV accuracy by 34.4% in video understanding tasks. Our SparseVLM can provide practical benefits for deploying off-the-shelf VLMs on edge devices and in the cloud setting.
Impact Statement
Our SparseVLM offers tangible advantages for implementing off-the-shelf VLMs on edge devices and in cloud environments. While our work carries no discernible societal implications, we see no necessity to underscore this aspect in the present context.
Appendix A Appendix
Figure 7: Comparison of visual redundancy in different vision tasks.
A.1 The Redundancy of Visual Tokens in VLM
In non-textual tasks like classification or segmentation, downsampling is commonly employed to decrease visual redundancy and enhance model training efficiency. Figure 7 illustrates this process, showing the reduction of tokens from 1166 to 576 in a downsampled image, resulting in a $50\%$ efficiency boost but a $15\%$ information loss (entropy decreased from 7.44 to 6.13). This trade-off is acceptable for such tasks. Conversely, for text-related tasks like visual question answering (VQA), which involve text and vision modalities, a distinct approach is needed. Highlighting the most information-dense text ($88\%$ of total text) alongside the region pertinent to the query in the image ($38\%$ of total image), we observe that image information is typically sparser than textual data. Hence, our SparseVLM method incrementally prunes visual token redundancy, maintaining crucial information for task accuracy. This strategy enhances model efficiency and effectiveness.
A.2 Compatibility with FlashAttention
To ensure compatibility between SparseVLM and FlashAttention (Dao et al., 2022) when extracting the matrix ${\bm{A}}$ or ${\bm{P}}$, we devise the dual-flash attention operation to directly obtain the average attention scores relative to the text raters. This operation is lightweight and enjoys the efficiency of FlashAttention. Specifically, the first forward pass operates identically to the original FlashAttention, generating the necessary hidden states. In the second forward pass, we introduce a specially designed ${\bm{V}}$ matrix. In this matrix, for the rows corresponding to the text raters we wish to analyze, we set the values to the reciprocal of the number of text raters. This configuration allows the inner product between the attention map and the ${\bm{V}}$ matrix to return the mean value of the attention scores for the selected text raters directly in FlashAttention. With the mean value, we perform a top-$k$ selection to identify the visual tokens to retain. Tokens that are excluded during this process are converted into masks, which are then applied to the hidden states produced by the first FlashAttention pass to complete the pruning operation. This method enables efficient integration of pruning with FlashAttention while preserving compatibility and computational efficiency. The specific principles and calculation of SparseVLM FlashAttention are as follows:
  1. 1.
    Attention Score Calculation. For each block $B$, compute the scaled dot-product attention scores as
    ${\bm{S}}_{B}=\frac{{\bm{Q}}_{B}{\bm{K}}_{B}^{T}}{\sqrt{d_{k}}},$
    where ${\bm{S}}_{B}$ is the attention score matrix computed within the block.
  2. 2.
    Block-wise Softmax. To ensure numerical stability, the Softmax is computed using the log-sum-exp trick as
    1. (a)
      Subtract the maximum value for numerical stability:
      ${\bm{S}}^{\prime}_{B}={\bm{S}}_{B}-\max({\bm{S}}_{B},\text{axis}=1)$
    2. (b)
      Normalize:
      ${\bm{P}}_{B}=\frac{\exp({\bm{S}}^{\prime}_{B})}{\sum\exp({\bm{S}}^{\prime}_{B},\text{axis}=1)}$
  3. 3.
    Special ${\bm{V}}$ Matrix. In order to return the mean value of the attention scores for the selected text raters directly with the FlashAttention, we need to design a special ${\bm{V}}$ matrix.
    ${\bm{V}}_{ij}=\begin{cases}1/n,&\text{if }i\in\{i_{1},i_{2},\dots,i_{k}\},\\ \\ 0,&\text{otherwise}.\end{cases}$
    Here, ${\bm{V}}$ is an $n\times d$ matrix, $n$ is the total number of rows in the matrix, $i$ is the row index, $1\leq i\leq n$, $\bm{s}=\{i\mid{\bm{r}}_{i}\geq m\},\,i\in\{1,2,\dots,L_{t}\}$ define the text raters which we selected in Section 3.2.
  4. 4.
    Incremental Accumulation. Rather than storing ${\bm{P}}$ explicitly, the result is directly accumulated into the output using:
    ${\bm{O}}_{B}={\bm{P}}_{B}\cdot{\bm{V}}_{B}$
    The final result is obtained by concatenating all blocks:
    ${\bm{O}}=\text{Concat}({\bm{O}}_{1},{\bm{O}}_{2},\ldots,{\bm{O}}_{B})$
  5. 5.
    Streaming Softmax. When combining multiple blocks, an incremental softmax computation ensures that normalization is maintained across the entire sequence:
    $\textrm{Softmax}({\bm{S}})=\exp({\bm{S}})/\sum\exp({\bm{S}})$
    This avoids global dependencies and enables efficient block-wise computation.
  6. 6.
    Top-$k$ Selection for Visual Tokens. The top-$k$ selection can be expressed as:
    ${\bm{O}}_{k}=\{x_{i}\in{\bm{O}}_{v}\mid\text{rank}(x_{i},{\bm{O}}_{v})\leq k\},$
    ${\bm{O}}_{v}=\{y_{j}\in\text{mean}({\bm{O}})\mid\text{visual tokens start}\leq j\leq\text{visual tokens end}\}.$
    where ${\bm{O}}=\textrm{Concat}({\bm{O}}_{1},{\bm{O}}_{2},\ldots,{\bm{O}}_{B})$ is the output array of the second FlashAttention, ${\bm{O}}_{v}$ is the visual tokens part of ${\bm{O}}$, $\text{rank}(x_{i},{\bm{O}}_{v})$ represents the position of $x_{i}$ in ${\bm{O}}_{v}$ when sorted in descending order.
    The corresponding indices of the top-$k$ elements are:
    ${\bm{I}}_{k}=\{i\mid x_{i}\in{\bm{O}}_{k}\}.$
  7. 7.
    Summary of SparseVLM with FlashAttention using Top-$k$ Selection. The complete process of SparseVLM FlashAttention can be summarized as
    ${\bm{I}}_{k}=\{i\mid x_{i}\in\{y_{j}\in{\bm{O}}_{v}\mid$ $\text{rank}(y_{j},\text{mean}(\textrm{Concat}\left(\bigcup_{B}\text{Softmax}\left(\frac{{\bm{Q}}_{B}{\bm{K}}_{B}^{T}}{\sqrt{d_{k}}}-\max({\bm{S}}_{B})\right)\cdot{\bm{V}}_{B}\right)$